How can i simplify (2^3-6^0)^-1
Answer by stanbon(75887)

How can i simplify (2^3-6^0)^-1
---
Work from the inside to the outside:
---
= (8-1)^-1
----
= 7^(-1)
---
= 1/7
=======================
Cheers,
Stan H.